OpenStudy (anonymous):

find the length: x=e^tcos t y=e^tsin t. t is betw 0 and 2

OpenStudy (anonymous):

hahha. if i were the real bieber do u think i would calll myself liliy and would i actually be smart enough to find this site and do math qs?

OpenStudy (anonymous):

to find the curve you do sqaure root of(dx/dy)^2+ (dy/dx)^2)

OpenStudy (anonymous):

ya?

OpenStudy (australopithecus):

are you looking for distance between y values?

OpenStudy (anonymous):

holy crap. DO nOT!!! come into MY QUESTION and write a whole long question like this. please open ur open question and i will be glad to help u out with these:)

OpenStudy (anonymous):

@liliy :o i did :P and no one answered it and theres only 4 questions but there not bath there career education .-.

OpenStudy (anonymous):

fine. im a nice person so ill help: but its really not fair what u did. A,A,A,D

OpenStudy (anonymous):

;p I'm sorry ill delete mine

OpenStudy (anonymous):

relax, its fine

OpenStudy (anonymous):

im just saying having it in ur own question is better.

OpenStudy (anonymous):

(also you can legit just google the questions)

OpenStudy (anonymous):

i did lmfao and I couldn't find any thing thats why I'm here

OpenStudy (anonymous):

alright:). well i hope i helped

OpenStudy (anonymous):

with what o.O?

OpenStudy (anonymous):

i just gave u the answers : 1.A 2.A 3.A 4.D

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

@liliy thanks :O

OpenStudy (dumbcow):

to find length of the parametric curve you need to integrate over the arc length of the curve \[\int\limits_{}^{}\sqrt{dx^{2} +dy^{2}}\] to factor in variable t, multiply by dt/dt \[\int\limits_{0}^{2}\sqrt{\frac{dx^{2} +dy^{2}}{dt^{2}}} dt\] which can be written as \[\int\limits_{0}^{2}\sqrt{(\frac{dx}{dt})^{2} +(\frac{dy}{dt})^{2}} dt\] \[\frac{dx}{dt} = e^{t}(\cos t -\sin t)\] \[\frac{dy}{dt} = e^{t} (\sin t + \cos t)\] \[\rightarrow \sqrt{2}\int\limits_{0}^{2}e^{t} dt\]

OpenStudy (anonymous):

wait, i dont undestand ur last step. wat happend to the sin and cos?

OpenStudy (dumbcow):

plug in the expressions for dx/dt and dy/dt , square them and combine terms when you do that you notice that sin^2 + cos^2 = 1 , and the rest of the sin and cos terms cancel so it simplifies to that nice easy integral i have above

OpenStudy (anonymous):

dx/dt is subtracting betw the cos and sin, so it doesnt cancel out!